A wall is built of two metals (stainless steel and copper) that are in thermal contact with each other. The temperature difference across the entire wall is 32.055 K and across the copper sheet it is 2.272 K. If the steel sheet is 1.163 cm thick, how thick is the copper sheet? cm
Answer #1

Using Fourier's law of heat conduction :  Q = \frac{k_{st} A \Delta T_{st}}{L_{st}} = \frac{k_{cu} A \Delta T_{cu}}{L_{cu}}

\Rightarrow \frac{k_{st} \Delta T_{st}}{L_{st}} = \frac{k_{cu} \Delta T_{cu}}{L_{cu}}

\Rightarrow L_{cu} = \frac{k_{cu} \Delta T_{cu}}{k_{st} \Delta T_{st}}L_{st}

\Rightarrow L_{cu} = \frac{385 *2.272}{50.2 * (32.055 - 2.272)}*1.163 = 0.68 cm\rightarrow Answer